OpenStudy (lime):

An airplane at T(80,20) needs to fly to both U(20,60) and V(110,85). What is the shortest possible distance for the trip?

OpenStudy (mathmate):

Compare the distances T-U-V and T-V-U and take the shorter of the two.

OpenStudy (lime):

\[\sqrt((20 - 60)^2 + (80 - 20)^2)\] \[\sqrt((-40)^2 + 60^2)\] \[\sqrt(1600 + 3600)\] \[\sqrt(5200)\]

OpenStudy (lime):

And then \[\sqrt((20 - 85)^2 + (80 - 110)^2)\]?

OpenStudy (mathmate):

Since the route T-U-V and T-V-U both use U-V, we know they are equal. So just compare which of T-U or T-V is shorter. one of them is sqrt(5200), and the other is sqrt(5125), so make your choice.

OpenStudy (anonymous):

And what type of information do you need to find the shortest distance? and how can you use a diagram to help you?